Circular No. 7796

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

PECULIAR VARIABLE IN MONOCEROS T. R. Geballe, Gemini Observatory; S. P. S. Eyres, University of Cen-

tral Lancashire; and A. Evans and V. H. Tyne, Keele University, report observations of this object (cf. IAUC 7785, 7786, 7791), made at the U.K. Infrared Telescope (+ CGS4) at ∼ Jan. 12.33 UT in nonphotometric conditions and at resolutions of 0.00125 and 0.0025 μm at 1.03–1.34 and 1.45–2.52 μm, respectively: “The spectrum shows a number of remarkable characteristics, including broad and shallow water bands at 1.4 and 1.9 μm; deep CO bands longward of 2.29 μm; hydrogen Brackett series lines in absorption, with the strongest absorptions occurring in the H band for transitions with n(upper) ∼ 14; and numerous strong lines of neutral metals (principally Mg, Si, and Fe) in the J band. Many of the metal lines have P-Cyg profiles, implying outflow velocities of several hundred km/s. The hydrogen lines do not show P-Cyg profiles. The continuum flux density decreases with wavelength and is consistent with the Rayleigh-Jeans falloff of a latetype star, with no indication of emission from hot dust. Rough infrared magnitudes are J = 7.0, H = 6.5, K = 6.3. The overall impression is of a cool, extended, and expanding atmosphere. Continued monitoring at infrared wavelengths is urged; higher spectral resolution is needed to resolve better the line profiles.”

S. Kimeswenger and C. Lederle, University of Innsbruck, report that 163 CCD frames were obtained with the Innsbruck 0.60-m telescope during Jan. 14.840–17.855 UT, showing no signature of a periodic variability in this variable star. Magnitudes (± 0.015): Jan. 14.886, B = 11.634, V = 9.802, R = 8.886, I_{c} = 7.998; 15.864, 11.733, 9.845, 8.935, 8.020; 17.838, 11.856, 9.955, 8.988, 8.057.

IM NORMAE G. J. Garradd, Loomberah, N.S.W., reports a precise position from an

unfiltered CCD image taken with a 0.45-m reflector on Jan. 17.715 UT: α = 15^{h}39^{m}26^{s}.465, δ = −52^{o}19′17′′.99 (equinox 2000.0; uncertainty ∼ 0′′.1 in each coordinate; ∼ 300 UCAC reference stars used), mag 10.5. Additional magnitude estimates (visual unless noted otherwise): 2001 Aug. 23, [14.5 (N. J. Brown, Quinns Rocks, W. Australia); Dec. 22.8, [10.4 (Brown; T-Max 400 film); 2002 Jan. 12.803, 8.3 (A. Pearce, Nedlands, W. Australia); 13.800, 8.3 (Pearce); 15.799, 8.2 (Pearce); 16.341, 7.66 (W. Liller, Viña del Mar, Chile; CCD broadband V; fog); 16.783, 8.7 (Pearce); 17.163, 8.9 (A. Amorim, Florianopolis, Brazil); 19.253, 8.9 (R. Shida, São Paulo, Brazil).

2002 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green